A nut store sells cashew for $4.00 per pound and peanuts for $1.50 per pound.
But at the end of the month , peanuts had not sold well so in order to sell
70 pounds of peanuts, the manager decided to mixt 70 pounds of peanuts with some cashews and sell the mixture for $3.50 per pound .
How many pounds of cashews should be mixed with the peanuts to ensure no change in the profit?
The manager should mix___pounds of c ashews with the peanuts.

Answer by Boreal(15235)

x=pounds of cashews
4x=revenue of cashew.
profit will be $105 (70 pounds of peanuts @$1.50 lb)
4x+105=(70+x)(3.50)
4x+105=245+3.50x
0.5x=140
x=280 pounds of cashews.
check
350 pounds @3.50 per pound is $1225.
280 pounds of cashews are $1120
70 pounds of peanuts are $105
That adds to $1225.
This makes sense, because the average price is going to be very close to what cashews sell for, so one would expect more cashews in the mixture.